OpenStudy (anonymous):

For each question, the rectangular coordinates of a point are given. Find the polar coordinates of each point. 4. (4, 0) 5. (3, 4) 6. (2, -2)

OpenStudy (anonymous):

(r, theta) can be found by r^2=x^2+y^2 tan(theta) = y/x

OpenStudy (anonymous):

let's do the first... r^2 = 4^2+0^2 ---> r^2 = 16 ---> r = 4 tan(theta) = 0/4 ---> tan(theta) = 0 ---> theta = 0 (4, 0) rectangular is (4, 0) in polar

OpenStudy (anonymous):

5. Polar ( r, θ ) = ( 5, 53°) r = √(3² + 4²) = 5 θ = arctan ( 4/3) = 53° Similarly, 6. ( 2√2, -45° )